OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (2, -5) with a slope of -3. y = -3x - 13 y = -3x + 11 y = -3x + 13 y = -3x + 1

OpenStudy (amistre64):

when given a set of options, we can easily determine a correct answer by plugging in the point values and see which one fits

OpenStudy (amistre64):

a point is defined as (x,y); in this case that gives us an x value of 2, and a y value of -5 plug in x=2 into the options to see what produces a -5

OpenStudy (anonymous):

y = -3x + 11

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@karenw12 @LaxPro @tornjeansxo @amber517

OpenStudy (anonymous):

did i do it

OpenStudy (anonymous):

y=mx+b y=(-3) · x+b y=(-3) · (2)+b (-5)=(-3) · (2)+b (-3) · (2)+b=-5 (-3)(2)+b=-5 (-6)+b=-5 b=1 y=-3x+1

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

np

OpenStudy (amistre64):

if y = -3x + 11 is correct, then when x=2, y=-5 y = -3(2) + 11 y = -6 + 11 y = 5 since y=5 is not the same as y=-5, this cannot be a correct option